OpenStudy (pucusana):

I need help. The volume of a right cylindrical water tank is 26,374.14 cubic cm. The area of the base of the cylinder is 720.63 square cm. What is the height (length) of the water tank? a) 109.796 cm b) 36.60 cm c) 27.449 cm d) 11.65 cm

OpenStudy (bobo-i-bo):

Volume of cylinder = area of base x height of cylinder

OpenStudy (pucusana):

I'm looking for the height.

OpenStudy (pucusana):

it has to be one of those options.

OpenStudy (bobo-i-bo):

So rearrange the formula I have given you to find the height, since you know what the volume and area of the cylinder is. :)

OpenStudy (pucusana):

my answer does not match any of the options.

OpenStudy (bobo-i-bo):

What is your answer, and how did you get the answer?

OpenStudy (jchick):

The radius is what of the base?

OpenStudy (pucusana):

my answer was 0.016

OpenStudy (bobo-i-bo):

How did you get your answer?

OpenStudy (pucusana):

26374.14/pi*720.63^2=0.016166

OpenStudy (bobo-i-bo):

720.63 is the AREA of the base of the cylinder. 720.63 is not the radius of the base of the cylinder

OpenStudy (jchick):

Oh wait sorry I misread the problem.

OpenStudy (jchick):

Yes @Bobo-i-bo that is correct the radius is half.

OpenStudy (pucusana):

oh

OpenStudy (bobo-i-bo):

You need to use my formula: volume of cylinder = area of base x height of cylinder

OpenStudy (pucusana):

but what's the half of 720.63?

OpenStudy (jchick):

Here is the formula for height.

OpenStudy (jchick):

Half of 720.63 is 360.315

OpenStudy (jchick):

You can find this by multiplying 720.63 by 0.5.

OpenStudy (bobo-i-bo):

I have no idea what you're doing @jchick ...

OpenStudy (jchick):

Finding the radius @Bobo-i-bo

OpenStudy (jchick):

All I did was find half of 720.63 @Bobo-i-bo

OpenStudy (pucusana):

so which means the answer is...

OpenStudy (bobo-i-bo):

why do you need to find the radius?

OpenStudy (pucusana):

I'm trying to find the height

OpenStudy (jchick):

We are trying to find the height of a right cylinder if I am not mistaken the formula asks for the radius.

OpenStudy (pucusana):

so the answer to the question is...

OpenStudy (jchick):

Do you want the answer or the method?

OpenStudy (pucusana):

the answer

OpenStudy (jchick):

I can't give that.

OpenStudy (mathmale):

No direct answers, please. Instead, show the work you have already done, so that we could give you appropriate feedback.

OpenStudy (pucusana):

i think 36.60 maybe?

OpenStudy (mathmale):

Volume of a vertical cylindrical tank: V=b*h, where b is the area of the base and h is the height. Please, show your calculations, as seeing them is the only way in which I can figure out your efforts.

OpenStudy (mathmale):

Let's say the volume is 26,374 cubic cm, and that the area of the base is 721 square cm. Using the formula I gave you, figure out the height. Hint: solve V=b*h for h.

OpenStudy (jchick):

@pucusana @mathmale is the best I know at math so you are in great hands with him!

OpenStudy (mathmale):

Thank you very much, @jchick!

OpenStudy (jchick):

Not a problem @mathmale because even when the problem confuses me you are right on track!

OpenStudy (mathmale):

Example: Supposing we had a right rectangular block of base area 100 square inches and volume 250 cubic inches. Since V = base * height, the height is (volume) / (base). Dividing the volume (250 cubic inches) by the base (100 square inches), we get 250 cu in -------- = 2.5 in (as the height of the block). Now, supposing that this right circular cyl. has a base area of 100 sq in

OpenStudy (mathmale):

721 square cm and that the volume is 26374 cubic cm. What is the height of the cylinder? Use precisely the same method as I did in the example, above.

OpenStudy (jchick):

@mathmale I have done some work I would appreciate if you could tell me if I am correct or not.

OpenStudy (jchick):

Like I said this problem was confusing me.

OpenStudy (mathmale):

Just divide the volume by the given area of the base. Even after rounding off the given info, the result (37 cm) is enough to enable us to find the correct answer (36.60 cm). Hope this problem now seems more solveable.

OpenStudy (jchick):

I came up with something different.

OpenStudy (mathmale):

@jchick : The work you shared looks like this:\[\frac{ A }{ 2 \pi r }=\frac{ 26374 }{ 2 \pi (360.31) }\]

OpenStudy (mathmale):

Recognizing that the height is the volume divided by the area of the base, my result is:\[\frac{ V }{ A }=\frac{ 26374 cm^3 }{ 721 cm^2 }=36.6 cm=height\]

OpenStudy (mathmale):

Note that your formula, 2pi(r), is for the circumference of the base, not for the area. Since the area is already known / given, and is 721 sq cm, just divide the volume by the area of the base to obtain the height of the cyl.

OpenStudy (jchick):

Ah ok I see my error.